OpenStudy (anonymous):

so is I wanted to make 1/6 and I didn't have 1/12 in measuring cup what would I have to do to get 1/6

OpenStudy (anonymous):

@AlexandervonHumboldt2 @pooja195 @Ghostgate @heyhihello12 @girlmeetsworld @Irwinlovesya @baller4life @babyboy2332 @Bee_see

OpenStudy (anonymous):

@Photon336 @mathmate @Miracrown @MTALHAHASSAN2 @Maharot

OpenStudy (mathmate):

Hint: what is \(\frac{1}{12}+\frac{1}{12}\)?

OpenStudy (anonymous):

1/6 but I have 2 measuring cups one has 1/4 2oz,1/3,1/2,2/3,3/4,1cup,1/4,1 1/2 that's one but idk if any of those equal 1/6 and the other is 1/2,2/3,3/4,1cup,1 1/4,1 1/3, 1 1/2, 1 2/3,1 3/4,2cup any of those numbers equal 1/6 @mathmate

OpenStudy (anonymous):

u there @heyhihello12

OpenStudy (heyhihello12):

What? what you need?

OpenStudy (heyhihello12):

mathmate is good he correct

OpenStudy (ghostgate):

Mathmate did get it, 1/12 + 1/12 = 1/6, however you want to get 1/6ths through your measuring cups current measurements, correct? As I mentioned in your previous problem where you originally had it, I would suggest to take about half of 1/3rds to get your 1/6ths measurement, because if you think about it 1/6 + 1/6 = 2/6 which is 1/3rds, so... out of just looking at it try your best to get half of 1/3rds. However, if you can find the equivalent to 1/12ths, then double that value and you get 1/6ths. Whichever works. I hope this helps in someway! Have a great day! ;) {---Ghostgate---}

OpenStudy (mathmate):

@alibaby You lost me when you said ``` 1/6 but I have 2 measuring cups one has 1/4 2oz,1/3,1/2,2/3,3/4,1cup, 1/4,1 1/2 that's one but idk if any of those equal 1/6 and the other is 1/2,2/3,3/4,1cup,1 1/4,1 1/3, 1 1/2, 1 2/3,1 3/4,2cup any of those numbers equal 1/6 @mathmate ``` Does it mean that you have 7 choices for the capacity of cup #1, and 10 choices for cup #2? and you want 1/6 oz or 1/6 cup? What is 2/3 - 1/2 cups?

OpenStudy (ghostgate):

Wow, 2/3 - 1/2 is something I didn't think of... nicely done mathmate. Lol - There's a reason you have a score of '99', right? xD

OpenStudy (mathmate):

Physically, if we have a 2/3 cup and a 1/2 cup, we fill the 2/3 cup and then empty into the 1/2 cup. 1/6 cup is what's left in the 2/3 cup.

OpenStudy (ghostgate):

I know, I just didn't think of that solution.